OpenStudy (anonymous):

3√x3y4z5 (the numbers are all powers)

OpenStudy (anonymous):

\[3\sqrt{x^{3}y^{4}z^{5}}\] ?

OpenStudy (anonymous):

yeah that's how it's supposed to look i could figure that out on my laptop

OpenStudy (anonymous):

maybe \[\sqrt[3]{x^3y^4z^5}\]

OpenStudy (anonymous):

\[3x y z \sqrt{x^{}y^{2}z^{3}}\]

OpenStudy (anonymous):

unless it is the cubed root, in which case it is \[xyz\sqrt[3]{yz^2}\]

OpenStudy (anonymous):

actually if it is the square root you can take out more but i am guessing it is the cubed root

OpenStudy (anonymous):

its the cubed root thank you for the help!!